# Emotional Strain and Frustration Drivers in LLM-Human Interactions in Software Engineering

### v1

@@ V-B Impact of frustrating experiences on motivation
Participant 39: “Sometimes the LLM hallucinates and puts me in loops. When I realize this, I resolve it myself using my human knowledge and 15 years of experience in the development field.”.
@@
Although frustration may not directly impact motivation, it remains a critical factor to consider due to its known connection with burnout. Sustained frustration is still significant as it contributes to emotional strain [tawfik2021frustration] —a known precursor to burnout in high-demand professions like software engineering. Thus, understanding and managing frustration, even when the motivation appears unaffected, is essential in supporting the long-term well-being of software engineers.
@@
Takeaway: Although frustration can be present in software engineers when LLMs fail to meet expectations, it generally does not diminish their motivation to complete tasks.
V-C Towards a less-frustrating user experience
@@
Several studies focus on enhancing LLMs; however, it is essential to recognise that these systems will likely always have room for improvement. Therefore, our focus in this study shifts to the human element. We propose enabling users with the knowledge and skills to navigate LLMs effectively to mitigate frustration and improve overall user experience. The goal is to address frustration triggers that arise during their use. Specifically, triggers such as misunderstanding of the intent, unsatisfying personal preferences, or even getting incorrect answers can be minimised by prompt engineering the query before sending it to the LLM. Prompt engineering can involve using prompt techniques (such as Few-shot learning), including relevant contextual information (e.g., description of the system) or constraints about the output (e.g., the output structure). Other frustration triggers, such as hallucinations and limitations of LLMs, can be mitigated if the engineers know the LLM’s capabilities and limitations and align their expectations. Since unmet expectations are among the primary frustration triggers[levine2022unhappy] (see Section IV-D), providing software engineers with training on effective usage and clear understanding of LLM capabilities to set realistic expectations can help reduce disappointment and enhance user satisfaction.
@@
Additionally, raising awareness about potential frustration triggers is important; engineers can manage their reactions accordingly if they recognise the likelihood of frustration in certain situations. For instance, using coping strategies rather than repeatedly attempting to elicit a perfect answer from the LLM. Therefore, to address the previous points, we suggest that software engineers (both in academia and industry) need training on how to use LLMs safely. This idea was also discussed by Barman et al. [barman2024beyond] where they propose providing guidelines for LLM users to know how to interact with different LLMs, for instance, if it is appropriate to generate artefacts or only to get some guidance. Our participants commented that they were very familiar with LLMs; however, most reported frustration, raising questions about whether they truly understand how to leverage LLMs effectively. Familiarity does not necessarily equate to proficiency[ladwig2012perceived], stressing the need for improved training and guidance on optimal usage strategies. A complete understanding of LLM capabilities and limitations can help users to manage their expectations.
Takeaway: Training on effective prompt engineering and a clear understanding of LLM capabilities and limitations to set realistic goals are important to address frustration triggers.
V-D Threats to Validity
@@
In this section we explain how we addressed this study’s threats to validity.
@@
External Validity: To address external validity and avoid overrepresenting certain groups or regions, we targeted respondents from different countries and several fields within software engineering. Similarly, we employed stratified sampling to ensure a balanced representation across demographics, skill levels, and industries.
@@
VI Conclusion
This study focused on the emotional strains, particularly frustration, experienced by software engineers when interacting with Large Language Models (LLMs) and not being assisted as intended. By identifying main triggers, such as the correctness and reliability of responses and issues related to personalisation, we emphasise that understanding the emotional impacts of LLM use in software engineering is important. This study’s insights bring attention to the potential risks to productivity and mental health if emotional responses go unaddressed. Future research should further explore the psychological implications of LLM use, focusing on sustainable strategies to support the well-being of software engineers and optimise their user experience with AI tools.

### v2

@@ 5.1. Frustrations in the Context of LLM Interaction
skipabove=1pt, innerlinewidth=1pt, innerlinecolor=customcolor, linewidth=0pt, backgroundcolor=customcolor!20 {mdframed} Takeaway: Frustration triggers studied in software engineer literature (including ours) come from spending extra time refining output, unlike in existing studies in other domains where it is due to performance and usability issues.
5.2. Impact of Frustrating Experiences on Motivation
While most of our participants expressed frustration (or similar emotions), their motivation to complete the task was not necessarily affected. This can suggest that although frustration may reside in the emotion wheel’s ’anger’ or ’irritation’ sections, the engineers’ resilience and coping mechanisms allowed them to manage these emotions without diminishing motivation.
Our results showed that participants often felt demotivated when an LLM failed to meet their expectations, such as when it did not assist them as intended, frequently leading to frustration. However, most of those who felt frustrated reported that their motivation remained intact, likely due to perceiving the LLM as merely one tool among others in achieving their goal. When an LLM could not provide the necessary assistance, participants commonly resorted to alternative solutions, such as using search engines like Google or relying on their expertise to complete the task independently. These observations align with findings by Franca et al. (França et al., 2014), who explored the connection between motivation and the satisfaction and happiness of software engineers, revealing that happiness slightly overlaps with but does not correlate with motivation.
@@ 5.2. Impact of Frustrating Experiences on Motivation
These findings suggest that, despite facing emotional challenges, software engineers maintain their motivation to persist with demanding tasks. Although frustration may not directly impact motivation, it remains a critical factor to consider due to its known connection with burnout. Sustained frustration is still significant as it contributes to emotional strain (Tawfik et al., 2021) —a known precursor to burnout in high-demand professions like SE. Thus, understanding and managing frustration, even when the motivation appears unaffected, is essential in supporting the long-term well-being of software engineers.
[…]
@@
skipabove=1pt, innerlinewidth=1pt, innerlinecolor=customcolor, linewidth=0pt, backgroundcolor=customcolor!20 {mdframed} Takeaway: Although frustration occurs when LLMs fail to meet expectations, it generally does not diminish the motivation to complete tasks.
5.3. Towards a Less-Frustrating User Experience
When designing chatbots and LLMs, it is essential to prioritise not only high accuracy and performance but also emotional intelligence such as recognising user frustration. Wilcox’ emotion wheel provides a nuanced view of emotional states, categorising emotions into primary and secondary feelings. Recognising and responding to these emotions in real-time is key.
@@ 5.3. Towards a Less-Frustrating User Experience
The emotional intelligence of LLMs has been explored by Wang et al. (Wang et al., 2023) where they saw that newer-generation LLMs such as GPT-4 (at the time of the study) show a better ability to understand user emotions that compared to humans’ emotional intelligence. However, recognising emotions is insufficient as the LLM should also know how to act accordingly. In Section 4.5, we saw that even minor LLM behaviours such as apologising after receiving feedback was making the engineers’ experience more frustrating. This was also described by Erlenhov et al (Erlenhov et al., 2019) about ideal development bots adapting communication to different individuals. While several studies focus on enhancing LLMs, we recognise that these systems will always have room for improvement. Therefore, our focus in this study shifts to the human element. We propose enabling users with the knowledge and skills to navigate LLMs effectively to reduce frustration and improve overall user experience. The goal is to address frustration triggers that arise during their use. Specifically, triggers such as misunderstanding of the intent, unsatisfying personal preferences, or even getting incorrect answers can be minimised by prompt engineering the query before sending it to the LLM. Prompt engineering can involve incorporating prompt techniques (such as Few-shot learning), relevant contextual information (e.g., system description), or constraints about the output (e.g., the output structure). Other frustration triggers, such as hallucinations and limitations of LLMs, can be minimised if the engineers use the LLM according to its capabilities and limitations. Since unmet expectations are among the primary frustration triggers(Levine et al., 2022) (see Section 4.4), providing software engineers with training on effective usage and clear understanding of LLM capabilities to set realistic expectations can help reduce disappointment and enhance user satisfaction.
@@
Additionally, raising awareness about potential frustration triggers is important; engineers can manage their reactions accordingly if they recognise the likelihood of frustration in certain situations. For instance, using coping strategies rather than repeatedly attempting to elicit a perfect answer from the LLM. Therefore, we suggest that software engineers need training on how to use LLMs safely. This idea was also discussed by Barman et al. (Barman et al., 2024) where they propose providing guidelines for LLM users to know how to interact with different LLMs, for instance, if it is appropriate to generate artefacts or only to get some guidance. Our participants commented that they were familiar with LLMs; however, most reported frustration, raising questions about whether they truly understand how to leverage LLMs effectively. Familiarity does not necessarily equate to proficiency(Ladwig et al., 2012), stressing the need for improved training and guidance on optimal usage strategies. A complete understanding of LLM capabilities and limitations can help users to manage their expectations.
[…]
skipabove=1pt, innerlinewidth=1pt, innerlinecolor=customcolor, linewidth=0pt, backgroundcolor=customcolor!20 {mdframed} Takeaway: A less frustrating experience arises from combining “emotionally intelligent” LLMs with engineers’ awareness to manage their expectations and reactions to frustration triggers.
5.4. Threats to Validity
@@
In this section, we explain the strategies to address this study’s threats to validity.
@@
External Validity: Our sample size of software practitioners and academics can limit to extent to which our findings can be generalized to the broader population of software engineers, to minimize this threat while avoiding overrepresenting certain groups or regions, we targeted respondents from different countries and several fields within SE. Similarly, we employed stratified sampling to ensure a balanced representation across demographics, skill levels, and industries.
@@
6. Conclusion
This study focused on the emotional strains, particularly frustration, experienced by software engineers when interacting with LLMs and not being assisted as intended. By identifying main triggers, such as the correctness and reliability of responses and issues related to personalisation, we emphasise that understanding the emotional impacts of LLM use in SE is important. This study’s insights bring attention to the potential risks to productivity and mental health if emotional responses go unaddressed. Future research should further explore the psychological implications of LLM use, focusing on sustainable strategies to support the well-being of software engineers and optimise their user experience with AI tools.